OpenStudy (anonymous):

A line passes through (1, -4) and has a slope of 3. What is an equation of the line in point-slope form?

OpenStudy (campbell_st):

the slope intercept of of a line is y = mx + b you know m = 3 so y = 3x + b substitute your point x = 1 and y = -4 into the equation and solve for b

OpenStudy (anonymous):

Ok so -4 = 3(1) =B?

OpenStudy (campbell_st):

that will give you the answer

OpenStudy (australopithecus):

https://www.youtube.com/watch?v=-6Fu2T_RSGM how to write it in point slope form

OpenStudy (anonymous):

Omg this is so confusing

OpenStudy (australopithecus):

Just google conversion from standard form (y=mx +b) to point slope form